OpenStudy (anonymous):

The following function defines a recursive sequence: f(0) = -3 f(1) = 6 f(n) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -3, 6, -9, -12, … -3, 20, -95, 480, … -3, 6, -9, 12, … -3, -20, -95, -480, …

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Ashleyrh1999

OpenStudy (anonymous):

you have: \(f(2)=-2*f(1)-f(0) = -2*6 - (-3) = -9\)

OpenStudy (anonymous):

now make for n = 3 and find out the answer

OpenStudy (anonymous):

c